OpenStudy (zuha):

5(2x+3) = 3 (4x+1)-2(3x+2) What is the value of x?

OpenStudy (anonymous):

10x+15=12x+3-6x-4 -2x+15=+3-6x-4 4x+15=-6x-1 4x=-6x+14 10x=14 whats 14divided by 10?

OpenStudy (zuha):

1.4?

OpenStudy (anonymous):

yes so x=1.4

OpenStudy (zuha):

Thank you :)

OpenStudy (anonymous):

you're welcome thanks for the medal